OpenStudy (anonymous):

A five-sided solid has the numbers 1, 2, 3, 4, and 5. What is the probability of rolling two five-sided number solids and getting a sum of either a 3 or an 8? A. 6/625 B. 9/625 C. 6/25 D. 1/5

OpenStudy (sidsiddhartha):

ok the five sided solid will have total 25 outcomes

OpenStudy (sidsiddhartha):

and u can have sum of three(3) by --(2+1)=3 and(1+2)=3 ok

OpenStudy (sidsiddhartha):

so we can have sum of three by 2 ways

OpenStudy (sidsiddhartha):

and we can have sum of 8 by-- (3+5)=8 (5+3)=8 (4+4)=8 so total 3 ways right?

OpenStudy (sidsiddhartha):

so getting a sum of either a 3 or an 8 is possible in=(3+2)=5 ways so your required probability \[P=5/25=1/5\]

OpenStudy (sidsiddhartha):

got it?

OpenStudy (anonymous):

yes thank you very much

OpenStudy (sidsiddhartha):

yw!!